Hi everyone my wife needs some help for her paper in finance i know is simple but i havent doen this for 15 years and i dont want to dissapoint her...
the formulas she gave me are the following ones..
K1=Rf+B(Km-Rf) this is the matrix formula now the variables..
B=1.25
Rf=?
Km=0.14
K1=0.15
second question variables:
B=1.10
Rf0.09
Km=?
K1=0.16
thank you for your help

Answer by stanbon(75887)

K1=Rf+B(Km-Rf) this is the matrix formula now the variables..
B=1.25
Rf=?
Km=0.14
K1=0.15
---------------
K1 = Rf + B*Km - B*Rf
---
Rf(1-B) = K1-BKm
--
Rf = (K1-BKm)/(1-B)
====
Rf = (0.15-1.25*0.14))/(1-1.25)
= -0.25/(-0.25)
= 1
======================================================
K1=Rf+B(Km-Rf)
second question variables:
B=1.10
Rf0.09
Km=?
K1=0.16
----
(Km-Rf) = (K1-Rf)/B
---
Km = [(K1-Rf)/B]+ Rf
-------
Km = [(0.16-0.09)/1.1] + 0.09
Km = (0.07)/(1.1) + 0.09
Km = 0.1536
=================
Cheers,
Stan H.